OpenStudy (countrygirl1431):

Evaluate 15 + |-3| - 4 8 14 22

OpenStudy (countrygirl1431):

@rishavraj

OpenStudy (welshfella):

+(- = -

rishavraj (rishavraj):

l-al = a

rishavraj (rishavraj):

@welshfella i believe thts modulous

OpenStudy (welshfella):

oh sorry that an absolute value I misread that

OpenStudy (welshfella):

I nee new specs lol!

rishavraj (rishavraj):

for sure u really need them....lol :P

OpenStudy (countrygirl1431):

lol well can some one help me

rishavraj (rishavraj):

yeah i jst gave u the hint :P l-al = a

OpenStudy (countrygirl1431):

lol im still confused

rishavraj (rishavraj):

hmmm u remember absolute value function http://www.purplemath.com/modules/solveabs.htm

OpenStudy (countrygirl1431):

yeah i know that

rishavraj (rishavraj):

so ??? l-3l = 3

OpenStudy (anonymous):

14

rishavraj (rishavraj):

@imdestroyer69 no direct answers -_-

OpenStudy (anonymous):

sorry xD i mean frog is the answer

OpenStudy (anonymous):

Task 2: School Fundraising Your school is planning a fundraising dinner. The expense for this event must not exceed $2,475.00. The team organizing the event has calculated that the cost per adult guest will be $18.00 and the cost per child guest will be $9.00. The venue can hold no more than 150 guests. a. Write a system of inequalities that represents this situation

OpenStudy (anonymous):

help plz